Write the equation you would use in excel to find the present value of the following cash flows assuming the required return was 10%/year and the long term perpetual growth rate in cash inflows post 68 months is 3%/year:

A B C D E F G
1 Time 0 8 Months 20 Months 32 Months 44 Months 56 Months 68 Months
2 -$200 m $40 m $80 m $120 m $140 m $180 m
Answer #1

Monthly ROR=10%/12=0.833%

Monthly Growth Rate=3%/12=0.25%

Terminal Cash Flow=180*(1+0.25%)/(0.833%-0.25%)=30951.97

So the Present Value of all the cash flow=

-200/(1+0.833%)^8 + 40/(1+0.833%)^20 + 80/(1+0.833%)^32 + 120/(1+0.833%)^44 + 140/(1+0.833%)^56 + 180/(1+0.833%)^68 + 30951.97/(1+0.833%)^69=$17644.11